Number of natural solutions of the equation l x + 3 l + l x - 5 l = 8

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Critical points= -3, 5
Consider the below cases in order to solve:
Case 1 : x less than -3
              -x-3-x+5=8
              x= -3 (not valid since x must be less than -3 in the given case)
 
Case 2 : -3 greater than equal to x less than 5
               x+3-x+5=8
               8=8
               Which is always true. Hence the whole case is valid. Natural no. values of x between -3 and 5 are 1, 2, 3, 4
 
 
Case 3: x greater than equal to 5
            x+3+x-5=8
            x=5
 
therefore natural no. values of x that satisfy the given equation are x= 1, 2, 3, 4, 5
5 values.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)